Phil has 30 ounces of a 60% sulfuric acid solution. How many ounces of 15% sulfuric acid solution should he add to obtain a solution that is 40% sulfuric acid?
Answer by Paul(988)

0.6(30)+0.15x=0.4(x+30)
18+0.15x=0.4x+12
-0.25x=-6
x=24

Hence about 24 ounces should be added to obtain 40% solution.
Paul.